OpenStudy (anonymous):

f(x)=-x^2-x-4 Determine a. if f has a maximum value or a minimum b. the value of x at which the max or min occurs c. the max or min value of f

OpenStudy (anonymous):

just graph it

myininaya (myininaya):

the parabola is oven downward so we have an upside down U that means we have a max

myininaya (myininaya):

open*

OpenStudy (anonymous):

a. maximum b. x = -1/2 c. -9/2

OpenStudy (anonymous):

A) derive, set equal to zero. There will be two. Test both critical points. If it goes from positive to negative, it's a max. If it's opposite, it's a min. If there is no change, which there won't be in this instance, it tells you nothing. B) the values you tested in a. C) plug the values in from b and solve for y.

myininaya (myininaya):

f(x)=-(x^2+x)-4=-(x^2+x+(1/2)^2)-4+(1/2)^2 =-(x+1/2)^2-4+1/4=(x+1/2)^2-15/4 so the vertex=(-1/2,-15/4)

OpenStudy (anonymous):

Correction, theres one critical point.